Circular No. 8733

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006du, 2006dv, 2006dw Three additional apparent supernovae have been discovered on unfil-

tered CCD images: 2006du by P. Luckas, O. Trondal, and M. Schwartz (0.35-m telescope, Perth; cf. IAUC 8713), 2006dv by T. Puckett and M. Peoples (0.50-m reflector, Ellijay, GA; cf. IAUC 8716), and 2006dw by N. Lee and W. Li (via LOSS/KAIT; cf. IAUC 8732).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006du July 18.9  0^{h}05^{m}12.^{s}26 −11^{o}29′23′′.6 17.0 13′′.2 W, 45′′.6 N
2006dv July 20.32 23 15 53.06 + 5 07 50.2 17.7 1′′.3 W, 6′′.9 N
2006dw July 23.25 16 17 43.29 +34 57 55.1 15.8 3′′.5 W, 1′′.7 N

Additional magnitudes for 2006du in IC 1529: 1983 Oct. 1, [21.0 (Digitized Sky Survey); 2006 June 10.8 UT, [18.0 (Luckas); July 22.9, 16.7 (Luckas). Additional magnitudes for 2006dv in UGC 12461: 2005 Aug. 9, [19.4; 2006 July 24.31, 17.5. Additional magnitudes for 2006dw: July 12.24, [19.5; 16.26, 18.3. The nucleus of the apparent host galaxy of 2006dw is located 37′′ east and 142′′ south of the center of NGC 6109.

COMET C/2006 M4 (SWAN) Twenty-six co-added 15-s CCD images taken by N. Teamo and S. F.

Hönig (0.35-m reflector, Punaauia, Tahiti) on July 17.2 UT show an 8′′ nuclear condensation surrounded by a fainter 45′′ coma (total mag 12.7) and a 2′.1 tail in p.a. 160^{o}. Three combined 30-s exposures by A. C. Gilmore (1.0-m reflector, Mt. John) on July 17.3 show a 2′.4 tail in p.a. 165^{o} and a well-condensed coma ∼ 40′′ in diameter.

Improved parabolic orbital elements to those published on IAUC 8729, taken from MPEC 2006-O17:

T = 2006 Sept.28.613 TT  ω = 62.^{o}851
Ω = 148.746 }2000.0
q = 0.77860 AU  i = 111.822

COMET 177P/BARNARD Visual total-magnitude estimates: July 7.07 UT, 13.2 (J. J. Gonzalez,

Leon, Spain, 0.20-m reflector); 16.42, 10.3 (D. A. J. Seargent, The Entrance, N.S.W., 25×100 binoculars); 19.91, 9.8 (M. Meyer, Frauenstein, Germany, 20×100 binoc.); 20.89, 9.1 (Meyer); 22.99, 8.5 (Gonzalez, 7×50 binoc.).

2006 July 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green